OpenStudy (kanard):

Simplify the following expression: (−5)(−3)( 4) −65 −60 60 65

OpenStudy (kanard):

@ganeshie8

OpenStudy (pawanyadav):

Simply multiply them As(-5)*(-3)=15 (15)* 4=?

OpenStudy (anonymous):

60

OpenStudy (pawanyadav):

Yes correct